A particle moves along the x axis and its position as a function of time is x(t) = (19.1 m/s)t - (1.9 m/s2)t2.
1
Expert's answer
2021-09-23T08:31:07-0400

"x=19.1t-1.9t^2,"

"v_{av}=\frac{\Delta x}{\Delta t}=\frac{x(t_2)-x(t_1)}{t_2-t_1},"

if "t_1=2.3~s," "t_2=3.9~s,"

"v_{av}=7.3~\frac ms."